starbucksgirl13:

For which function will both ordered pairs (3, 6) and (4, 8) be solutions? y = x + 4 y = 12 - x y = 2 x y = x + 3

loganisbored:

umm

loganisbored:

so whats 3x6?

starbucksgirl13:

24

loganisbored:

nope

Trios:

Don’t you need to use an equation for this?

starbucksgirl13:

18

loganisbored:

yes

loganisbored:

whats 4x8

starbucksgirl13:

@Timmyspu

Timmyspu:

What do you need help with?

starbucksgirl13:

this qusestion

loganisbored:

he has been summoned

loganisbored:

@dude

Timmyspu:

So what is the first thing we have to do?

jhonyy9:

For which function will both ordered pairs (3, 6) and (4, 8) be solutions? y = x + 4 y = 12 - x y = 2 x y = x + 3 so just you need to check for what function are roots these ordered pairs this you can take it than you substitute the x value in place of x and the y value in place of y ,make the calcule and you need to get that the left side is equal right side hope helped understandably easy

Timmyspu:

Good job @jhonyy9